OpenStudy (anonymous):

Please Help Me!! "Using complete sentences, describe the benefit of simplifying rational expressions before multiplying." I really just need the example though, I don't know how to do this!

OpenStudy (anonymous):

I am so sorry!, this is what i meant to ask! "Using complete sentences, describe how you would check the quotient of two rational expressions. Create and demonstrate your own example of this concept."

OpenStudy (anonymous):

It makes sure that when you do multiply, you don't have a more complicated situation to work with after doing so. By simplifying, you get variables and like terms removed and organized so that your next step is actually something that gets you to a solution.

OpenStudy (anonymous):

if you have the quotient, then multiply that by what was in the denominator and see if what you get is the numerator you started with. P(x)/Q(x)=S(x). A polynomial divided by another polynomial equals another one. Take S(x)*Q(x)=? P(x). This applies to any rational expression that does solve to be a quotient. Polynomials are just a common instance in algebra and what not.

OpenStudy (anonymous):

I'm not upto come up with a specific set right now, but that shouold illustrate the idea for checking your question.

OpenStudy (anonymous):

Thank you!